# The Adoption and Usage of AI Agents:

### v1

@@ 1 Introduction
We illustrate the hierarchical structure of our agentic taxonomy in Figure 1 and report our complete taxonomy in Table 1. Productivity is the dominant topic with a 36% share. It is followed by learning (21%), media (16%), and shopping (10%). The most prevalent subtopics with over 5% query share include courses (13%), goods shopping (9%), research (8%), document editing (8%), account management (7%), and social media (7%). The most frequently observed tasks are exercise assistance (9%), research information summarization and analysis (7%), document creation and editing (7%), product search and filtering (6%), and research information search and filtering (6%). We also study the use of agents across environments, which are the websites on which these tasks are performed. The concentration of environments varies substantially across subtopics: the top five environments account for 97% of queries in music, 97% in videos, and 96% in professional networking, compared to only 28% in account management, 35% in shopping for services, and 37% in project management. Across all use cases, 55% of agentic queries originate from personal use settings, 30% from professional use settings, and 16% from educational use settings. In the short term, users show strong within-topic persistence, demonstrating stickiness in use cases; when topic transitions occur, they are more likely to migrate toward productivity, learning, or media topics. Over time, query shares shift from travel and media topics to productivity, learning, and career topics.
@@ 2 Related Work
Our paper is directly related to the literature on how people use LLMs and AI agents in real-world settings. [^18] Following our definition of AI agents, we do not discuss papers that do not involve the agent taking actions to manipulate their environments. Our paper extends recent work on the adoption and usage of LLM chatbots. Several prominent studies have examined this topic, including handa2025economic, which documents user interactions with Claude, and zhao2024wildchat and chatterji2025people, which analyze how people use ChatGPT. Anthropic has also released detailed analyses focusing on specific user groups, such as university students (handa2025education), educators (benthand2025education), and different geographies and enterprises (appelmccrorytamkin2025geoapi). In addition, aubakirova2025stateofai uses OpenRouter data to study LLM chatbot usage across both open and closed-source models. These papers developed taxonomies to categorize standard Q&A queries. We also create a taxonomy using internal data from an AI product; however, our key contribution is the focus on agentic queries. The main difference is that Q&A queries focus on information exchange between the user and model in a conversation. In contrast, agentic queries focus on the agent executing tasks on behalf of the user in an external environment.
Evidence on how people use AI agents in the field is more limited and usually focused on specific use cases, such as coding. For example, claudecode studies the usage of Claude Code, a coding agent, in software development, and sarkar2025ai investigates the adoption, usage, and productivity impact of coding agents in Cursor. Our contribution differs in that we analyze a general-purpose agent operating across all common use cases. [^19] Although not directly related to the focus of our paper, it is worth noting adjacent research that examines the adoption and usage of LLM chatbots through user surveys (e.g., humlum2025unequal, bick2024rapid, handa2025interviewer); the productivity and performance impact of LLM chatbots across various occupations and tasks through field (e.g., dell2023navigating, wiles2024genai, brynjolfsson2025generative, cui2025effects, vendraminelli2025genai) and lab experiments (e.g., noy2023experimental, peng2023impact, merali2024scaling); and the behavior of AI agents and human–agent collaboration through case studies (e.g., anthropic2025vend), firm surveys (e.g., pan2025measuring), or lab experiments (e.g., allouah2025your, ju2025collaborating).
@@ 5.1 Adoption and Usage Intensity
Tables 8 and 9 in Appendix A show the top ten occupation subclusters and their user shares by agent adopter shares and AAR, and agentic query share and AUR, respectively. Patterns at the subcluster level are largely consistent with those at the cluster level, with software engineers being the largest subcluster, accounting for 14% of adopters and 15% of queries, and having AAR and AUR around 1.1 and 1.2, respectively. All other subclusters are below 6% in both adopter and query shares. Subclusters in marketing—such as business development and sales, digital marketing, and market research—and in entrepreneurship—such as information management, operations, and strategy—tend to have the highest AAR and AUR.
@@ 5.2 Use Cases
Tasks under topics and subtopics capture the low-level tasks the agent is expected to complete to achieve the end goals. We show the top ten tasks in Table 6. Half of the top ten are in learning, including various learning and research assistance. The other five are split across productivity (edit documents and manage account settings), shopping (search or summarize product information), and media (search social media).
@@
Table 17 in Appendix A shows the top five tasks in each occupation cluster. In general, research, document editing, and shopping-related tasks appear consistently across clusters. Some occupation clusters feature a prominent task. For instance, search products appear in 21% of queries from the advanced manufacturing cluster. Other clusters, in contrast, have a more diffuse task composition. The top tasks in the entrepreneurship and design clusters—summarizing research information and searching products—appear in fewer than 8% of their queries. The top tasks by occupation also shed light on why certain occupations tend to adopt and use the agent more. Knowledge-intensive sectors such as digital technology, entrepreneurship, finance, and academia tend to use the agent for research and learning-related tasks. In contrast, highly digitized sectors such as marketing and design tend to use the agent for media-related tasks.
@@
Environment refers to the external world with which the agent interacts while performing a task to achieve its goals. In our context, the environment is the specific website on which the agent operates for a given query. We show the top environments by overall query shares in Table 7. The top 10 environments together account for 63% of queries; they are typically the leading websites in their corresponding domains.
We break down the top five environments under each subtopic and their query shares in Tables 18, 19, 20, 21, 22, and 23 in Appendix A. A single environment dominates some subtopics. For instance, linkedin.com alone accounts for 93% of queries in professional networking, and the query shares of youtube.com and docs.google.com323232docs.google.com includes Google Docs, Sheets, Slides, and Forms.—the top environments in video and spreadsheet editing subtopics—are more than twenty times larger than the share of the second environments in those subtopics. On the other hand, there is only a 2% difference between coursera.org and netacad.com under courses, and a 3% difference between instagram.com and x.com under social media. Table 24 in Appendix A shows the combined shares of the top five environments in each subtopic. A higher share indicates the agent usage is more concentrated in a few environments. The level of concentration varies significantly: the top five environments account for 97% of queries in music, 97% in videos, and 96% in professional networking, compared to only 28% in account management, 35% in services shopping, and 37% in project management, respectively. Lastly, Table 25 in Appendix A shows the top 5 environments for each occupation cluster, which are closely related to the main use cases for each occupation.
@@ 6 Discussion
We note a few important caveats of our dataset. First, because Comet is a new product, our sample primarily reflects early adopters, who may skew toward more tech-savvy users. We characterize these early adopters using an internal survey in Appendix C. Relatedly, given the short time span of our data, we do not systematically investigate changes in usage patterns over time, and any longitudinal results should be interpreted within this context. Second, the classification of an agentic query depends on internal query understanding modules that trigger the agent based on predicted query intent. These intent predictors show high prediction accuracy in internal validation studies; nonetheless, the data may include both false positives (when a non-agentic intent triggers the agent) and false negatives (when an agentic intent does not trigger the agent). [^34] Two sets of classifiers determine agent activation: one for browser control and one for specific apps. The browser control classifier is a supervised model trained on user queries labeled as either showing or not showing agentic intent. Five features are used for prediction: the current query, the currently viewed page, the user’s previous queries in the same conversation, enabled connectors, and the number of attachments. The classifier achieves an ROC-AUC of 0.95 and both precision and recall of 0.90 at the optimal threshold. Each connector additionally has its own classifier, following a similar process. Similarly, the classification into our agentic taxonomy, usage context, and occupation clusters also contains noise. Third, although AI agents’ autonomy and task horizons continue to expand, our results should not be interpreted as suggesting any particular balance between automation and augmentation in use cases. For instance, an agentic session may appear to be automation, but users may be breaking a task into smaller pieces and delegating only some subtasks to the agent, which is closer to an augmentation case. A comprehensive treatment on such a topic would require having a complete picture of how users manage their workflows outside of Comet.
@@ Appendix D Validation of Agent Use Cases Classifier
We randomly select 1,000 agentic queries for manual labeling. Note that because of our focus on agentic queries, we cannot evaluate the classifier’s performance on public Q&A query datasets such as WildChat. [^37] https://wildchat.allen.ai/ These queries are sampled from a larger set previously classified as containing no harmful content. To maintain representativeness, we retain queries that may contain personal information and redact names, email addresses, addresses, and phone numbers. Each query in the sample is independently labeled by two or three annotators using our agentic taxonomy. Out of the 1,000 queries, 370 show disagreement among annotators on what the primary topic and subtopic are. Most disagreements stemmed from variations in labeling quality across annotators. For each query that shows disagreement, our team manually reviews it and labels it against our taxonomy. The final golden dataset includes the 630 queries on which all annotators agreed, along with the 370 queries we labeled. We validate the classifier against the golden dataset and across multiple runs, and the agreement rates are listed in Table 30.
@@
  • Note: The table shows the agreement rate between the classifier label and the golden dataset and across runs. The tasks are specific to subtopics, so when the classifier disagrees with the golden dataset or with itself across runs, the tasks will by definition have zero agreement rates. Therefore, the task agreement rate is conditional on topic- and subtopic-level agreement. The across-runs agreement rate is the average pairwise agreement rate across three runs.

### v2

@@ 1 Introduction
We illustrate the hierarchical structure of our agentic taxonomy in Figure 1 and report our complete taxonomy in Table 1. Productivity is the dominant topic with a 36% share. It is followed by learning (21%), media (16%), and shopping (10%). The most prevalent subtopics with over 5% query share include courses (13%), goods shopping (9%), research (8%), document editing (8%), account management (7%), and social media (7%). The most frequently observed tasks are exercise assistance (9%), research information summarization and analysis (7%), document creation and editing (7%), product search and filtering (6%), and research information search and filtering (6%). We also study the use of agents across environments, which are the websites on which these tasks are performed. The concentration of environments varies substantially across subtopics: the top 5 environments account for 97% of queries in music, 97% in videos, and 96% in professional networking, compared to only 28% in account management, 35% in shopping for services, and 37% in project management. Across all use cases, 55% of agentic queries originate from personal use settings, 30% from professional use settings, and 16% from educational use settings. In the short term, users show strong within-topic persistence, demonstrating stickiness in use cases; when topic transitions occur, they are more likely to migrate toward productivity, learning, or media topics. Over time, query shares shift from travel and media topics to productivity, learning, and career topics.
@@ 2 Related Work
Our paper is directly related to the literature on how people use LLMs and AI agents in real-world settings. [^18] Following our definition of AI agents, we do not discuss papers that do not involve the agent taking actions to manipulate their environments. Our paper extends recent work on the adoption and usage of LLM chatbots. Several prominent studies have examined this topic, including handa2025economic, which documents user interactions with Claude, and zhao2024wildchat and chatterji2025people, which analyze how people use ChatGPT. Anthropic has also released detailed analyses focusing on specific user groups, such as university students (handa2025education), educators (benthand2025education), and different geographies and enterprises (appelmccrorytamkin2025geoapi). In addition, aubakirova2025stateofai uses OpenRouter data to study LLM chatbot usage across both open- and closed-source models. These papers developed taxonomies to categorize standard Q&A queries. We also create a taxonomy using internal data from an AI product; however, our key contribution is the focus on agentic queries. The main difference is that Q&A queries focus on information exchange between the user and model in a conversation. In contrast, agentic queries focus on the agent executing tasks on the user’s behalf in an external environment.
Evidence on how people use AI agents in the field is limited and typically focuses on specific use cases, such as coding. For example, claudecode studies the usage of Claude Code, a coding agent, in software development, and sarkar2025ai investigates the adoption, usage, and productivity impact of coding agents in Cursor. Our contribution differs in that we analyze a general-purpose agent operating across all common use cases. [^19] Although not directly related to the focus of our paper, it is worth noting adjacent research that examines the adoption and usage of LLM chatbots through user surveys (e.g., humlum2025unequal, bick2024rapid, handa2025interviewer); the productivity and performance impact of LLM chatbots across various occupations and tasks through field (e.g., dell2023navigating, wiles2024genai, brynjolfsson2025generative, cui2025effects, vendraminelli2025genai) and lab experiments (e.g., noy2023experimental, peng2023impact, merali2024scaling); and the behavior of AI agents and human–agent collaboration through case studies (e.g., anthropic2025vend), firm surveys (e.g., pan2025measuring), or lab experiments (e.g., allouah2025your, ju2025collaborating).
@@ 5.1 Adoption and Usage Intensity
Tables 8 and 9 in Appendix A show the top 10 occupation subclusters and their user shares by agent adopter shares and AAR, and agentic query share and AUR, respectively. Patterns at the subcluster level are largely consistent with those at the cluster level, with software engineers being the largest subcluster, accounting for 14% of adopters and 15% of queries, and having AAR and AUR around 1.1 and 1.2, respectively. All other subclusters are below 6% in both adopter and query shares. Subclusters in marketing—such as business development and sales, digital marketing, and market research—and in entrepreneurship—such as information management, operations, and strategy—tend to have the highest AAR and AUR.
@@ 5.2 Use Cases
Tasks under topics and subtopics capture the low-level tasks the agent is expected to complete to achieve the end goals. We show the top 10 tasks in Table 6. Half of the top 10 are in learning, including various learning and research assistance. The other five are split across productivity (edit documents and manage account settings), shopping (search or summarize product information), and media (search social media).
@@
Table 17 in Appendix A shows the top 5 tasks in each occupation cluster. In general, research, document editing, and shopping-related tasks appear consistently across clusters. Some occupation clusters feature a prominent task. For instance, search products appear in 21% of queries from the advanced manufacturing cluster. Other clusters, in contrast, have a more diffuse task composition. The top tasks in the entrepreneurship and design clusters—summarizing research information and searching products—appear in fewer than 8% of their queries. The top tasks by occupation also shed light on why certain occupations tend to adopt and use the agent more. Knowledge-intensive sectors such as digital technology, entrepreneurship, finance, and academia tend to use the agent for research and learning-related tasks. In contrast, highly digitized sectors such as marketing and design tend to use the agent for media-related tasks.
@@
Environment refers to the external world with which the agent interacts while performing a task to achieve its goals. In our context, the environment is the specific website on which the agent operates for a given query. We show the top environments by overall query shares in Table 7. The top 16 environments together account for 64% of queries, with the top 5 alone representing 43%. These environments are typically the dominant websites within their respective domains.
We break down the top 5 environments under each subtopic and their query shares in Tables 18, 19, 20, 21, 22, and 23 in Appendix A. A single environment dominates some subtopics. For instance, linkedin.com alone accounts for 93% of queries in professional networking, and the query shares of youtube.com and docs.google.com323232docs.google.com includes Google Docs, Sheets, Slides, and Forms.—the top environments in video and spreadsheet editing subtopics—are more than twenty times larger than the share of the second environments in those subtopics. On the other hand, there is only a 2% difference between coursera.org and netacad.com under courses, and a 3% difference between instagram.com and x.com under social media. Table 24 in Appendix A shows the combined shares of the top 5 environments in each subtopic. A higher share indicates the agent usage is more concentrated in a few environments. The level of concentration varies significantly: the top 5 environments account for 97% of queries in music, 97% in videos, and 96% in professional networking, compared to only 28% in account management, 35% in services shopping, and 37% in project management, respectively. Lastly, Table 25 in Appendix A shows the top 5 environments for each occupation cluster, which are closely related to the main use cases for each occupation.
@@ 6 Discussion
We note a few important caveats of our dataset. First, because Comet is a new product, our sample primarily reflects early adopters, who may skew toward more tech-savvy users. We characterize these early adopters using an internal survey in Appendix C. Relatedly, given the short time span of our data, we do not systematically investigate changes in usage patterns over time, and any longitudinal results should be interpreted within this context. Second, the classification of an agentic query depends on internal query understanding modules that trigger the agent based on predicted query intent. These intent predictors show high prediction accuracy in internal validation studies; nonetheless, the data may include both false positives (when a non-agentic intent triggers the agent) and false negatives (when an agentic intent does not trigger the agent). [^34] Two sets of classifiers determine agent activation: one for browser control and one for specific apps. The browser control classifier is a supervised model trained on user queries labeled as either showing or not showing agentic intent. Five features are used for prediction: the current query, the currently viewed page, the user’s previous queries in the same conversation, enabled connectors, and the number of attachments. The classifier achieves an ROC-AUC of 0.95 and both precision and recall of 0.90 at the optimal threshold. Each connector additionally has its own classifier, following a similar process. Similarly, the classification into our agentic taxonomy, usage context, and occupation clusters also contains noise. Third, although AI agents’ autonomy and task horizons continue to expand, our results should not be interpreted as suggesting any particular balance between automation and augmentation in use cases. For instance, an agentic session may appear to be automation, but users may break a task into smaller pieces and delegate only some subtasks to the agent, which is closer to an augmentation case. A comprehensive treatment on such a topic would require having a complete picture of how users manage their workflows outside of Comet.
@@ Appendix D Validation of Agent Use Cases Classifier
We randomly select 1,000 agentic queries for manual labeling. [^37] Note that because of our focus on agentic queries, we cannot evaluate the classifier’s performance on public Q&A query datasets such as WildChat. https://wildchat.allen.ai/ These queries are sampled from a larger set previously classified as containing no harmful content. To preserve representativeness, we retain queries that may include personal information, while redacting names, email addresses, physical addresses, and phone numbers. Each query in the sample is independently labeled by two or three annotators using our agentic taxonomy. Out of the 1,000 queries, 370 show disagreement among annotators on what the primary topic and subtopic are. Most disagreements stemmed from variations in labeling quality across annotators. For each query that shows disagreement, our team manually reviews it and labels it against our taxonomy. The final golden dataset includes the 630 queries on which all annotators agreed, along with the 370 queries we labeled. We validate the classifier against the golden dataset and across multiple runs, and the agreement rates are listed in Table 30.
@@
  • Note: The table shows the agreement rate between the classifier label and the golden dataset and across runs. The tasks are specific to subtopics, so when the classifier disagrees with the golden dataset or across runs on topics or subtopics, the tasks will by definition have zero agreement rates. Therefore, the task agreement rate is conditional on topic- and subtopic-level agreement. The across-runs agreement rate is the average pairwise agreement rate across three runs.